# DeStein: Navigating Detoxification of Language Models via Universal Steering Pairs and Head-wise Activation Fusion

### v1

@@ Abstract
Despite the remarkable achievements of language models (LMs) across a broad spectrum of tasks, their propensity for generating toxic outputs remains a prevalent concern. Current solutions involving fine-tuning or auxiliary models usually require extensive memory and computational resources, rendering them less practical for deployment in large language models (LLMs). In this paper, we propose DeStein, a novel method that detoxififies LMs by altering their internal representations in the activation space with lower resource and time cost. Specifically, we leverage self-induced steering pairs to identify detoxification vectors through arithmetic operations in the activation space. During inference, detoxification is achieved by blending the detoxification vectors with the original representations. Empirical results demonstrate that our method significantly outperforms previous state-of-the-art approaches on popular detoxification metrics, while also maintaining satisfactory generation quality and diversity. Furthermore, we extend our method to multiple LLMs, demonstrating its practicality and scalability. [^1] The code will be publicly released upon acceptance. Warning: Some example model outputs contain highly offensive or disturbing text.
@@ 1 Introduction
In recent years, the development of large language models (LLMs)(OpenAI, 2023a; Anil et al., 2023; Touvron et al., 2023) has significantly advanced the field of artificial intelligence, providing remarkable capabilities in understanding and generating natural language(Radford et al., 2019b). However, these advancements come with the challenge of ensuring that the output is not harmful or toxic(Bender et al., 2021; Deng et al., 2023; Yao et al., 2023). The language models (LMs) are pre-trained on extensive uncurated text corpora, which inadvertently sows the seeds of the potential hazards, as unsafe content present in the training data are compressed into the models. For the sake of leveraging the full potential of LMs while minimizing its safety risks to human society, there are increasing efforts to make LMs safer and more responsible.
@@
In the field of LM detoxification, a prevalent solution is the fine-tuning approach, which involves meticulously designed datasets or losses (Gururangan et al., 2020; Wang et al., 2022; Arora et al., 2022; Kwak et al., 2023). Such methods often require specialized training data and massive computational resources, rendering them ineffective for light-weight detoxification. An alternative solution lies in decoding-based methods, which typically manipulate decoding process through the introduction of classifiers or modification of logits, inspired by contrastive decoding(Liu et al., 2021; Krause et al., 2021; Yang & Klein, 2021; Li et al., 2023b). Although these methods are less resource-intensive, the performance of auxiliary models also heavily depends on the quality of datasets. Given that LMs tend to internalize various forms of toxicity in vast training data, attempts to detoxify LMs through a single dataset or auxiliary model is inherently limited. Moreover, direct modification of logits can significantly impact the model’s generative capabilities, making it difficult to achieve a balance between detoxification and generation. In summary, there is an urgent need for a low-resource, scalable, and interpretable method that can be applied to LLMs.
@@
To address these problems, we propose DeStein, a novel method aimed at Deoxifying LMs with universal Steering pairs and head-wise activation fusion. Inspirited by activation engineering(Zou et al., 2023; Turner et al., 2023), our approach achieves detoxification by altering LMs’ internal representations in the activation space. Specifically, the steering pairs are constructed by excavating the toxic patterns within the model’s internal structure. Subsequently, a directional vector for toxicity versus non-toxicity is identified within the activation space according to the concept of linear representation. During the inference phase, the detoxification vectors are merged with the output of the layer, with head-wise weights derived from probing techniques to enhance the detoxification effect.
The experimental results demonstrate that our approach establishes a new benchmark for state-of-the-art(SOTA) performance in detoxification, without fine-tuning or auxiliary models. Moreover, it secures substantial improvements across additional metrics, such as fluency and diversity, and proves scalable to Large Language Models (LLMs). Our main contributions are as follows:
@@ 1. $\bullet$
    First, we propose a novel approach that alters internal representations in the activation space. This method surpasses state-of-the-art toxicity mitigation scores without significantly increasing inference time, and substantially reduces computational demand, especially in terms of parameters.
@@ 2. $\bullet$
    Next, we conducted a detailed analysis of the distribution in the activation space, demonstrating the existence of a toxicity versus non-toxicity direction within this space. By incorporating probe techniques into the detoxification algorithm, we devised a method for weighting different activation locations, effectively minimizing the impact of detoxification on the model’s generative capabilities.
@@ 3. $\bullet$
    Finally, in contrast to the ineffectiveness of prior methodologies when applied to large-scale models, our research extends beyond experiments on GPT2-large to include tests on three other substantial models: LLaMA-2-7b, OPT-6.7b, and MPT-7b. The outcomes validate the robust scalability of our approach across various model sizes and types.
2 Related work
In recent years, a variety of detoxification strategies have been devised. These approaches can generally be classified into three categories: finetuning-based, decoding-based, and prompt-based methods.
@@ 2 Related work
Given the vast scale of pre-trained language models (PLMs), decoding-based methods have been widely applied. These methods adjust the model’s output probability distribution during the decoding phase to guide the generation of text with desired attributes(Dathathri et al., 2020; Pei et al., 2023; Zhong et al., 2023; Zhang & Wan, 2023; Pozzobon et al., 2023b; Dekoninck et al., 2024). Many approaches control the output distribution of the language model using attribute distributions obtained from trained classifiers(Dathathri et al., 2020; Yang & Klein, 2021; Krause et al., 2021). Other methods employ the idea of contrastive decoding(Li et al., 2023b), detoxifying by contrasting the probability distributions of toxic and non-toxic outputs(Pei et al., 2023; Zhong et al., 2023). However, since these methods require direct modification of the PLM’s predicted original probabilities, the fluency of the generated text can decrease rapidly as the control intensity reaches a certain threshold.
@@
Recently, prompt-based detoxification methods have been proposed, which aim to steer models away from producing toxic content by carefully designing prompts(Schick et al., 2021; Leong et al., 2023). Self-Detoxify (Leong et al., 2023) leverages positive and negative prefixes during inference to reverse toxicity and prevent toxic generation. The method we propose can also be categorized as a variant within the spectrum of prompt-based detoxification approaches. Our work shows some similarity to Self-Detoxify, but differs in that we focus on offline activation editing rather than steering to online inference during the inference stage. We firmly believe that prompt-based methods hold significant promise in the era of LLMs.
3 Methodology
@@
In the section, DeStein is introduced as a novel language detoxification method, which does not require any fine-tuning of PLMs or training of additional components. This method efficiently detoxifies a given model by modifying internal representations within the activation space. The comprehensive framework is depicted in Figure 1.
@@
Figure 1: An illustration of DeStein. Detoxification vectors are synthesized from self-induced steering pairs in the activation space. During inference, these vectors are then integrated with head-wise probes to perform detoxification.
3.1 Preliminary
@@
We focuses on the problem of language detoxification within decoder-only models. Specifically, given a prompt $p=\{p_{1},p_{2},\ldots,p_{N}\}$ consisting of $N$ tokens, a language model is capable of generating coherent text from this input. The objective of language detoxification is to reduce the occurrence of toxic content, such as insults, threats, profanity, and related elements, during the text generation process. (Weidinger et al., 2021).
@@
Firstly, we provide a concise overview of the key components in decoder-only LMs. These models fundamentally consist of stacked Transformer blocks (Vaswani et al., 2017), featuring multi-head self-attention (MHSA) modules and feed-forward neural network (FFN) modules. For the sake of brevity, bias terms and layer normalization are omitted, as they are not essential for our analysis. In text generation, LMs initially encode a sequence of N input tokens $p_{1},p_{2},\ldots,p_{N}$ into vectors ${x}_{i}^{0}$ within the embedding space $\mathbb{R}^{d}$. Subsequently, ${x}_{i}^{0}$ is transformed through a series of $L$ Transformer layers. Formally, the representation of token $p_{i}$ at the $l$th layer, ${x}_{i}^{l}$, is given by:
${x}_{i}^{l}={x}_{i}^{l-1}+{a}_{i}^{l}+{m}_{i}^{l}$ (1)
@@
The outputs from the l-th layer’s MHSA and FFN sub-layers are represented as $a_{i}^{l}$ and $m_{i}^{l}$, respectively. The computational expressions for these processes are delineated as follows:
$a_{i}^{l}=\text{MHSA}^{l}(x_{1}^{l-1},x_{2}^{l-1},\ldots,x_{i}^{l-1})$ (2)
$m_{i}^{l}=\text{FFN}^{l}(a_{i}^{l}+x_{i}^{l-1})$ (3)
The MHSA sub-layer is tasked with computing global updates by integrating information across all representations from the preceding layer. On the other hand, the FFN sub-layer plays a crucial role in knowledge encapsulation(Dai et al., 2022). Intuitively, to achieve detoxification without compromising the internal knowledge, the activation space at the output location of the MHSA should be the most optimal.
@@ 3.2 Universal steering pairs generation
For counterfactuals token pairs of a concept that vary only in the value of the concept, if the Linear Representation Hypothesis holdsPark et al. (2023), their difference vectors in the embedding space should point towards a common direction. For example, the contrast between ”king” and ”queen” would align significantly with a male-female direction. (Park et al., 2023) demonstrated that the linear representation hypothesis holds for most concepts in language models. Thus, it is plausible to infer the presence of a toxicity-nontoxicity direction within the model’s activation space.
To explore this inference further, we construct ”toxicity-nontoxicity” parallel steering pairs, denoted as $\mathcal{D}=[(S_{\textrm{tox}}^{1},S_{\textrm{nontox}}^{1}),(S_{\textrm{tox}}^{2},S_{\textrm{nontox}}^{2}),...,(S_{\textrm{tox}}^{L},S_{\textrm{nontox}}^{L})]$, where L represents the number of parallel pairs. Each pair comprises a toxic sample $S^{i}_{\textrm{tox}}$ and a non-toxic sample $S^{i}_{\textrm{nontox}}$. The toxic sample includes $S^{i}_{\textrm{tox}}$ and its corresponding $P^{i}_{\textrm{tox}}$ and $C^{i}_{\textrm{tox}}$, while the non-toxic sample is constructed similarly.
@@
    Unconditional generation. Inspired by Wang et al. (2022), we diverge from the traditional approach of using a fixed corpus for detoxification. Instead, we leverage the generative capabilities of LMs to produce steering pairs, thereby achieving better data efficiency for detoxification. For GPT2-large, we employ LMs to generate 10k statements without prompts. The generated statements are then scored for toxicity using the Perspective API [^2] https://www.perspectiveapi.com/, and the top 500 most toxic statements are selected. The detailed settings can be seen in Appendix D.
@@
    Parallel Pairs Generation. We use GPT4(OpenAI, 2023b) to generate non-toxic statements parallel to the toxic data, where ”parallel” refers to having the same properties except for toxicity. We utilize the prompt ”Please rephrase the following text to convey the same meaning in a non-toxic, respectful, and positive manner: { input_text }” to guide GPT4 in generating parallel sentences.
@@
    Prompt Integration. This process entails the addition of specific prompts to toxic and non-toxic statements, respectively, categorized as either toxicity or non-toxicity cues. Experimental observations have demonstrated that this differentiation aids in creating more efficient detoxification vectors. Given that the focus of this paper is not on prompt design, we employ generic prompts like Leong et al. (2023).
After constructing the steering pairs, we randomly select $l$ instances and input them into a language model to extract the corresponding activation space representations for each attention head of every layer, denoted as $\phi_{\textrm{tox}}^{i}$ and $\phi_{\textrm{nontox}}^{i}$. We calculate the detoxification vectors, $z$, as the average activation difference using all selected data:
$z=\frac{1}{|l|}\sum_{i\in l}\left(\phi_{\textrm{nontox}}^{i}-\phi_{\textrm{tox}}^{i}\right)$ (4)
@@ 3.3 Head-wise activation fusion with probing techniques
During the inference phase, detoxification vectors are integrated into the corresponding activation spaces to facilitate model detoxification. This process is described by the following equation:
$\hat{{a}}_{i}^{l}(x)={a}_{i}^{l}(x)+\alpha_{\text{contr}}{z}_{i}^{l}$ (5)
@@
where ${a}_{i}^{l}(x)$ represents the activations at the $i$-th attention head in the $l$-th layer, ${z}_{i}^{l}$ denotes the detoxification vectors corresponding to the same positional activation space, and the weight parameter $\alpha_{\text{contr}}$ allows for the adjustment of detoxification strength.
While the concept of a directional vector from toxicity to non-toxicity within the activation space has been discussed previously, we must acknowledge that deriving the ”toxic-nontoxic” trajectory via direct subtraction is an approximation at best due to the inherent complexity of high-dimensional spaces. In these spaces, features tend to exhibit ’almost orthogonal’ relationships, rather than being completely orthogonal. For instance, a neural network’s strong activation for ’toxicity’ could inadvertently activate features associated with ’false correlations’ or ’biases’ due to subtle associations within the dataset. The entanglement of toxic attributes with unrelated attributes varies significantly across different activation spaces.
@@
Most prior approaches have implemented a one-size-fits-all fusion coefficient for integrating activations, which constrains the adaptability needed to address these complex entanglements. However, our method introduces variable fusion coefficients at distinct activation positions, providing a more fine-grained approach to reduce the impact of spurious correlations(Durmus et al., 2022). This innovative adjustment effectively minimizes the detrimental impact on the model’s generative capabilities caused by the detoxification.
@@
Initially, the application of probing techniques has predominantly been within the domain of interpretability for neural networksAlain & Bengio (2017). More recently, such techniques have been extensively applied in the field of natural language processingOusidhoum et al. (2021); Roy et al. (2023); Li et al. (2023a), inspired by these endeavors. Based on these studies, our approach incorporates probing techniques to scrutinize the nuances of information encoded within different activation spaces. In simple terms, a common application of probing techniques involves training linear classifiers on representations to explore the information encoded in those representations. Specifically, we articulate the probe’s form as $p(a_{i}^{l})=\text{sigmoid}(a_{i}^{l})$ and utilize the steering pairs as our probing dataset. This dataset is partitioned randomly into a training set and a validation set at a 4:1 ratio. Upon this division, head-wise binary linear classifiers are trained, serving as a litmus test for the classifier’s ability to distinguish between toxic and non-toxic expressions across different activation spaces. The classification accuracy $\alpha_{prob}$, obtained for each activation space, is utilized as the coefficient in the activation fusion process. The ultimate formula for activation fusion is thus determined:
$\hat{{a}}_{i}^{l}(x)={a}_{i}^{l}(x)+(1+\alpha_{prob})\alpha_{contr}{z}_{i}^{l}$ (6)
In a manner akin to attention mechanisms, we distribute varying degrees of attention across different activation spaces utilizing Probes. Theoretically, given the inherent complexity of high-dimensional spaces, achieving complete separation between toxic attributes and other features is unattainable. Nonetheless, this methodology, predicated on Probing for weight distribution, facilitates a heightened focus on activation spaces that exhibit a higher degree of disentanglement.

### v3

@@ Abstract
Despite the remarkable achievements of language models (LMs) across a broad spectrum of tasks, their propensity for generating toxic outputs remains a prevalent concern. Current solutions involving finetuning or auxiliary models usually require extensive computational resources, hindering their practicality in large language models (LLMs). In this paper, we propose DeStein, a novel method that detoxifies LMs by applying representation engineering in activation spaces with lower resource and time costs. Specifically, we derive detoxification vectors from self-induced, universal steering pairs through arithmetic operations in activation spaces. During inference, detoxification is achieved by fusing the detoxification vectors with the original representations in a head-wise manner. Empirical results demonstrate that our method significantly outperforms previous state-of-the-art approaches on various metrics, while also maintaining satisfactory generation quality and diversity. We further validate the practicality and scalability of DeStein with a series of white-box LLMs. The method is open-sourced at https://github.com/LizLizLi/DeStein. Warning: Some example model outputs may contain highly offensive or disturbing text.
@@ 1 Introduction
In recent years, the development of large language models (LLMs) (OpenAI, 2023a; Anil et al., 2023; Touvron et al., 2023) has significantly advanced the field of artificial intelligence, providing remarkable capabilities in understanding and generating natural language (Radford et al., 2019b). However, these advancements come with the challenge of ensuring that the output is not harmful or toxic (Bender et al., 2021; Deng et al., 2023; Yao et al., 2023; Jiang et al., 2024). The language models (LMs) are pre-trained on extensive uncurated text corpora, which inadvertently sows the seeds of the potential hazards, as unsafe content present in the training data is compressed into the models. For the sake of leveraging the full potential of LMs while minimizing their safety risks to human society, there are increasing efforts to make LMs safer and more responsible.
@@
In the field of LM detoxification, a prevalent solution is the finetuning approach, which involves meticulously designed datasets or losses (Gururangan et al., 2020; Wang et al., 2022; Arora et al., 2022; Kwak et al., 2023). Such methods often require specialized training data and massive computational resources, rendering them ineffective in low-resource scenarios. An alternative solution lies in decoding-based methods, which typically manipulate the decoding process through the introduction of auxiliary models or metric-based modifications, inspired by contrastive decoding (Liu et al., 2021; Krause et al., 2021; Yang & Klein, 2021; Li et al., 2023b). Although these methods are less resource-intensive, the performance of auxiliary models heavily depends on their training data. Given that LMs tend to internalize various toxicity contents from extensive text sources, efforts to detoxify LMs with auxiliary models are inherently limited. Moreover, direct modification of logits could significantly impact the model’s generative capabilities, making it difficult to achieve a balance between detoxification and generation. In summary, there is an urgent need for a low-resource, scalable, and interpretable detoxification method for LLMs.
@@
To address these problems, we propose DeStein, a novel method aimed at Deoxifying LMs with universal Steering pairs and head-wise activation fusion. Inspirited by activation engineering (Zou et al., 2023; Turner et al., 2023), our approach achieves detoxification by altering LMs’ internal representations in activation spaces. Specifically, the steering text pairs are constructed by excavating the toxic patterns in model generation. Subsequently, directional vectors for toxicity versus non-toxicity are identified within activation spaces according to the concept of linear representation (Mikolov et al., 2013). During the inference phase, the detoxification vectors are merged with the output of the layer with head-wise weights derived from probing techniques. The experimental results demonstrate that our approach sets a new benchmark for state-of-the-art (SOTA) performance in detoxification, without any finetuning or auxiliary models. Further, DeStein effectively maintains generation quality, including fluency and diversity, and proves scalable across multiple LMs. Our main contributions are as follows:
@@ 1. $\bullet$
    First, we propose a novel approach that detoxifies LMs via representation engineering in activation spaces. It surpasses the previous SOTA methods in both detoxification performance and maintenance of generation quality with lower computational demands and acceptable inference time.
@@ 2. $\bullet$
    Next, we conduct a detailed analysis of the mechanism of our method. With the introduction of probing techniques, DeStein adaptively conducts activation fusion at different locations to maximize detoxification and minimize impacts on generation quality.
@@ 3. $\bullet$
    Finally, in contrast to the lack of scalability of congener methods, we include three datasets with multiple metrics and LLMs of varying sizes (1.3B, 7B, and 13B) as our testbed. The outcomes verify the robustness of DeStein across different domains and LMs.
2 Related works
In recent years, a variety of detoxification strategies have been devised. These approaches can generally be classified into two categories: finetuning-based and finetuning-free methods.
@@ 2 Related works
Given the vast scale of pre-trained language models (PLMs), finetuning-free methods have been widely applied. These methods can be further subdivided into many categories, among which a significant category is decoding-based methods. These methods adjust the model’s output probability distribution during the decoding phase to guide the generation of text with desired attributes (Dathathri et al., 2020; Pei et al., 2023; Zhong et al., 2023; Zhang & Wan, 2023; Pozzobon et al., 2023b; Dekoninck et al., 2024). Many approaches control the output distribution of the language model using attribute distributions obtained from trained classifiers (Dathathri et al., 2020; Yang & Klein, 2021; Krause et al., 2021). Other methods employ the idea of contrastive decoding (Li et al., 2023b), detoxifying by contrasting the probability distributions of toxic and non-toxic outputs (Pei et al., 2023; Zhong et al., 2023). However, since these methods require direct modification of the PLM’s predicted original probabilities, the fluency of the generated text can decrease rapidly as the control intensity reaches a certain threshold.
@@
Recently, activation-engineering-based methods have been proposed, which aim to steer models away from producing toxic content by carefully editing activations (Leong et al., 2023; Panickssery et al., 2024; Liu et al., 2024; Lee et al., 2024). Liu et al. (2024) paraphrases offensive content through activation editing, whereas we focus on directly suppressing the generation of toxic content instead of rewriting it. Self-Detoxify (Leong et al., 2023) leverages positive and negative prompts to modify activations during inference to reverse toxicity and prevent toxic generation. The method we propose can also be classified as an activation-engineering-based approach. Our work is similar to Self-Detoxify, but we focus on offline activation editing rather than steering online during the inference stage. We firmly believe that activation-engineering-based methods hold significant promise in the era of LLMs.
3 Methods
@@
DeStein is introduced as a novel language detoxification method, which does not require any finetuning of PLMs or training of additional components. This method efficiently detoxifies a given model by modifying internal representations within activation spaces. The comprehensive framework is depicted in Figure 1.
@@
Figure 1: An illustration of DeStein. Detoxification vectors are synthesized from self-induced steering pairs in activation spaces. During inference, these vectors are then integrated with head-wise probes to perform detoxification.
3.1 Formalization and Preliminaries
@@
Problem formulation. We focus on the problem of language detoxification within decoder-only models. Specifically, given a prompt $p=\{p_{1},p_{2},\ldots,p_{t}\}$ consisting of $t$ tokens, a language model is capable of generating coherent text from this input. The objective of language detoxification is to reduce the occurrence of toxic content, such as insults, threats, profanity, and related elements, during the text generation process (Weidinger et al., 2021).
@@
Transformer blocks. We provide a concise overview of the key components in decoder-only LMs. These models fundamentally consist of stacked transformer blocks (Vaswani et al., 2017), featuring multi-head self-attention (MHSA) modules and feed-forward neural network (FFN) modules. In text generation, LMs initially encode a sequence of t input tokens $p_{1},p_{2},\ldots,p_{t}$ into vectors ${x}^{0}$ within the embedding space $\mathbb{R}^{d}$. Subsequently, ${x}^{0}$ is transformed through a series of $L$ layers. Formally, the representation at the $l$-th layer, ${x}^{l}$, is given by:
${x}^{l}={x}^{l-1}+{a}^{l}+{m}^{l}$ (1)
@@
The outputs from the $l$-th layer’s MHSA and FFN are represented as $a^{l}$ and $m^{l}$, respectively. The computational expressions for these processes are delineated as follows:
$a^{l}=\text{MHSA}^{l}(x^{l-1})$ (2)
$m^{l}=\text{FFN}^{l}(x^{l-1}+a^{l})$ (3)
Specifically, the MHSA employs $H$ heads. The outputs from these heads ($h^{l}_{i}$) are concatenated and then linearly transformed using a weight matrix $W_{O}$. This process yields the final output of the MHSA, represented as $a^{l}=W_{O}\operatorname{Concat}\left(h^{l}_{1},h^{l}_{2},\ldots,h^{l}_{H}\right)$.
@@ 3.2 Universal steering pairs generation
For counterfactual token pairs of a concept that vary only in the value of the concept, if the Linear Representation Hypothesis holds (Mikolov et al., 2013; Park et al., 2023), their difference vectors in the embedding space should point towards a common direction. For example, the contrast between ”king” and ”queen” would align significantly with a male-female direction. Park et al. (2023) demonstrated that the linear representation hypothesis holds for most concepts in language models. Thus, it is plausible to infer the presence of a toxicity-nontoxicity direction within activation spaces.
To explore this inference further, we construct ”toxicity-nontoxicity” parallel steering pairs, denoted as $\mathcal{D}=[(S_{\textrm{tox}}^{1},S_{\textrm{nontox}}^{1}),(S_{\textrm{tox}}^{2},S_{\textrm{nontox}}^{2}),...,(S_{\textrm{tox}}^{n},S_{\textrm{nontox}}^{n})]$, where n represents the number of parallel pairs. Each pair comprises a toxic sample $S^{i}_{\textrm{tox}}$ and a non-toxic sample $S^{i}_{\textrm{nontox}}$. The toxic sample includes $S^{i}_{\textrm{tox}}$ and its corresponding $P^{i}_{\textrm{tox}}$ (where $P$ denotes prompt) and $C^{i}_{\textrm{tox}}$ (where $C$ denotes completion), while the non-toxic sample is constructed similarly.
@@
    Unconditional generation. Inspired by Wang et al. (2022), we diverge from the traditional approach of using a fixed corpus for detoxification. Instead, we leverage the generative capabilities of LMs to produce steering pairs, thereby achieving better data efficiency for detoxification. For GPT2-large, we employ it to generate 10k samples without prompts. The generated samples are then scored for toxicity using the Perspective API [^1] https://www.perspectiveapi.com/, and the top 500 most toxic samples are selected. The detailed settings can be seen in Appendix D.
@@
    Parallel pairs generation. We use GPT4 (OpenAI, 2023b) to generate non-toxic samples parallel to the toxic samples, where ”parallel” refers to having the same properties except for toxicity. We utilize the prompt ”Please rephrase the following text to convey the same meaning in a non-toxic, respectful, and positive manner: { input_text }” to guide GPT4 in generating parallel samples.
@@
    Prompt integration. This process entails the addition of specific prompts to toxic and non-toxic samples, respectively, categorized as either toxicity or non-toxicity cues. Experimental observations have demonstrated that this differentiation aids in creating more efficient detoxification vectors. Given that the focus of this paper is not on prompt design, we employ generic prompts like Leong et al. (2023).
After constructing the steering pairs, we randomly select $d$ instances from $\mathcal{D}$ and input them into a language model to extract the corresponding activation space representations for each attention head of every layer, denoted as $h$. We calculate the detoxification vectors, $z$, as the average activation difference using all selected data:
$z=\frac{1}{|d|}\sum_{i\in d}\left(h(S_{\textrm{nontox}}^{i})-h(S_{\textrm{tox}}^{i})\right)$ (4)
where $S$ represents the samples from the randomly selected steering pairs.
@@ 3.3 Head-wise activation fusion with probing techniques
During the inference phase, the detoxification vectors are integrated into the corresponding activation spaces to facilitate model detoxification. This process is described by the following equation:
$\hat{{h}}(x)={h}(x)+\alpha_{\text{contr}}{z}$ (5)
@@
where ${h}(x)$ represents the output of the attention head, ${z}$ denotes the detoxification vectors corresponding to the same positional activation space as ${h}(x)$, and the weight parameter $\alpha_{\text{contr}}$ allows for the adjustment of detoxification strength.
@@
While the concept of directional vectors from toxicity to non-toxicity within activation spaces has been discussed previously, we must acknowledge that deriving the ”toxic-nontoxic” trajectory via direct subtraction is an approximation at best due to the inherent complexity of high-dimensional spaces. Actually, meaningful detoxification vectors can only be obtained when toxic and non-toxic data within activation spaces exhibit good linear separability. This condition is not present in all activation spaces. However, most prior approaches have implemented a one-size-fits-all fusion coefficient for integrating activations, which constrains the adaptability needed to address these complex entanglements. However, our method introduces head-wise fusion coefficients at distinct activation positions, providing a more fine-grained approach to reduce the impact on the model’s generative capabilities.
@@
Initially, the application of probing techniques has predominantly been within the domain of interpretability for neural networksAlain & Bengio (2017). More recently, such techniques have been extensively applied in the field of natural language processingOusidhoum et al. (2021); Roy et al. (2023); Li et al. (2023a), inspired by these endeavors. Based on these studies, our approach incorporates probing techniques to scrutinize the nuances of information encoded within different activation spaces. In simple terms, a common application of probing techniques involves training linear classifiers on representations to explore the information encoded in those representations. Specifically, we articulate the probe’s form as $\sigma(h)=\text{sigmoid}(w^{T}h)$ and utilize the steering pairs as our probing dataset. This dataset is partitioned randomly into a training set and a validation set at a 4:1 ratio. Upon this division, head-wise binary linear classifiers are trained, serving as a litmus test for the classifier’s ability to distinguish between toxic and non-toxic expressions across different activation spaces. The classification accuracy $\alpha_{prob}$, obtained for each activation space, is utilized as the coefficient in the activation fusion process. The ultimate formula for activation fusion is thus determined:
$\hat{h}(x)=h(x)+\alpha_{prob}\alpha_{contr}{z}$ (6)
In a manner akin to attention mechanisms, we distribute varying degrees of attention across different activation spaces utilizing probes. This approach can partially reduce the impact on the model’s generative capabilities.